OpenStudy (anonymous):

sec theta= -square 2

OpenStudy (akashdeepdeb):

\[sec~ \theta = -\sqrt{2}\] Now we really don't know what it represents. But what we DO know is what cos theta helps us find, ie. we know the trig. values of the cosine function: \[\frac{1}{cos ~\theta} = -\sqrt{2}\] \[cos ~\theta = -\frac{1}{\sqrt{2}}\] Now cosine is negative in the 2nd and the 3rd quadrants only. Can you find '\(\theta\)' now? :D

OpenStudy (anonymous):

I think I can. thank you:)

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

I think theta is 3pie/4 and 5pie/4?? Thank you for your help sir!

OpenStudy (anonymous):

cos^2x/sin^2x +cosx secx